A saline solution contains 3% salt. How much water should be added to 100 mililiters of this solution to dilute it to a 2% solution?
Thank you

Answer by mananth(16946)

---------percent ---------------- quantity
solution 3%---------------- 100
water 0%---------------- x
Total 2%---------------- 100+x
...
3*100+0 *x=2(100 +x)
300=200+ 2x
-2x=200 -300
-2x=-100
/-2
x=50 milliliter